Solve the rational equation:
1/x-1/(x+1)=1/56
Thank you in advance for any help.

Answer by scott8148(6628)

multiply by [56x(x+1)] (the LCD) ___ 56x + 56 - 56x = x^2 + x

x^2 + x - 56 = 0

factor the quadratic to find x